OpenStudy (anonymous):

Solve each system using Elimination: 2x+4y=22 2x-2y=-8. how do you subtract the equations from each other?

OpenStudy (anonymous):

This one is easy. You can get rid of the 2x just by multiplying one of the equations by -1. Let's do it to the second one. So now we have this 2x+4y=22 -2x+2y=8

OpenStudy (the_fizicx99):

He gave you the system set out, now you just have to add them, the 2x and -2x cancel out.

OpenStudy (anonymous):

I don't think she's gonna do it. It was 12 minutes ago I put that!